OpenStudy (anonymous):

Integral cosz dz/(z (z-2) (z-4)) at |z|=1

terenzreignz (terenzreignz):

\[\Large \oint\limits_\gamma^{}\frac{\cos(z)}{z(z-1)(z-4)}dz\qquad \gamma = \{z \ \ | \ \ |z|=1\}\]

terenzreignz (terenzreignz):

Looks daunting? :D

terenzreignz (terenzreignz):

@rrakes Stay with me ^_^

OpenStudy (anonymous):

i need solution and the concept to solve.

terenzreignz (terenzreignz):

Gladly. First, Cauchy-Goursat theorem. This integrand is analytic everywhere except for a few points known as poles. Can you identify the poles?

OpenStudy (anonymous):

yes 0 2 4

terenzreignz (terenzreignz):

2? I don't think so :)

OpenStudy (anonymous):

z-2 is there in denominator and at z=2 the integrand is not defined

terenzreignz (terenzreignz):

There is no z-2 in your post -_- No matter: Is this the correct integral: \[\Large \oint\limits_\gamma^{}\frac{\cos(z)}{z(z-2)(z-4)}dz\]

terenzreignz (terenzreignz):

@rrakes This is kind of important ^_^

OpenStudy (anonymous):

yes